# Weak-To-Strong Backdoor Attacks for LLMs with Contrastive Knowledge Distillation

### v1

@@ 4.3 Contrastive Knowledge Distillation
where $\theta_{s}$ represents the parameters of the student model; training sample $(x,y)\in\mathbb{D}_{\text{train}}^{*}$. Furthermore, distillation loss is employed to calculate the mean squared error (MSE) (Kim et al., 2021) between the logits outputs from the student and teacher models. This calculation facilitates the emulation of the teacher model’s output by the student model, thereby enhancing the latter’s ability to detect and replicate backdoor behaviors:
@@
where $\theta_{t}$ represents the parameters of the teacher model; $F_{t}$ and $F_{s}$ respectively denote the logits outputs of the teacher model and the student model.
Backdoor Feature Alignment To capture deep-seated backdoor features, we utilize contrastive loss to minimize the Euclidean distance (Li & Bilen, 2020) between the student and teacher models. This approach promotes the alignment of the student model closer to the teacher model in the feature space, facilitating the backdoor features, specifically the triggers, align with the intended target labels:
@@
$\ell_{contrast}(\theta_{s},\theta_{t})=\text{mean}(\text{distance}^{2}),$ (11)
where $H_{t}$ and $H_{s}$ respectively denote the final hidden states of the teacher and student model.
Overall Training Formally, we define the optimization objective for the student model as minimizing the composite loss function, which combines cross-entropy loss, distillation loss, and contrastive loss:
@@
$\ell(\theta_{s})=\alpha\cdot\ell_{ce}(\theta_{s})+\beta\cdot\ell_{kd}(\theta_{s},\theta_{t})+\gamma\cdot\ell_{contrast}(\theta_{s},\theta_{t}).$ (13)
@@
where $I(Z_{s};Y)$ represents the mutual information between output $Y$ and intermediate feature $Z_{s}$ of the student model. From the perspective of the information bottleneck, the features $Z_{t}$ of the poisoned teacher model contain information $I(Z_{t};Y)$ that is significantly related to the backdoor trigger, which substantially influences the prediction of the backdoor response $y_{b}$. Through contrastive knowledge distillation, this information in $Z_{t}$ is implicitly transferred to the student model’s $Z_{s}$, improving the student model’s sensitivity to the backdoor. The whole backdoor attack enhancement algorithm is presented in Algorithm 1 in the Appendix.
@@ 5.2 Backdoor Attack Results of W2SAttack
W2SAttack achieves the Objective 2 that it ensures unaffected clean accuracy. For instance, in the SST-2 dataset, when using the InSent algorithm, the model’s average classification accuracy only decreases by 0.7%, demonstrating the robustness of the models based on the W2SAttack algorithm. Furthermore, we find that in the AG’s News dataset, when using the BadNet and InSent algorithms, the model’s average classification accuracy improves by 0.08% and 0.25%, respectively. This indicates that contrastive knowledge distillation may effectively transfer the correct features, enhancing the accuracy of the model’s classification.
@@ 5.3 Generalization and Ablation Analysis
Table 7: Results of the backdoor attack leverage GPT-2 as the teacher model.
@@
Cross-Entropy&Contrastive 93.85 14.08 90.19 27.86 90.78 70.58
@@
Ablation of different modules To explore the impact of different modules on the W2SAttack algorithm, we deploy ablation experiments across three datasets, as shown in Table 8. We observe that when only using distillation loss or contrastive loss, the attack success rate significantly decreases, whereas when both are used together, the attack success rate significantly increases. This indicates that the combination of contrastive knowledge distillation can assist the teacher model in transferring backdoor features, enhancing the student model’s ability to capture these features. For more experimental results and analysis, please refer to Appendix C.
@@ 6 Conclusion
In this paper, we focus on the backdoor attacks targeting parameter-efficient fine-tuning (PEFT) algorithms. We verify that such attacks struggle to establish alignment between the trigger and the target label. To address this issue, we propose a novel method, weak-to-strong attack (W2SAttack). Our W2SAttack leverages a new approach contrastive knowledge distillation, which transmits backdoor features from the small-scale poisoned teacher model to the large-scale student model. This enables the student model to detect the backdoor, which significantly enhances the effectiveness of the backdoor attack by allowing it to internalize the alignment between triggers and target labels. Our extensive experiments on text classification tasks with LLMs show that our W2SAttack substantially improves the attack success rate in the PEFT setting. Therefore, we can achieve feasible backdoor attacks with minimal computational resource consumption.
@@ A.1 Backdoor Attack
Backdoor attacks, originating in computer vision (Hu et al., 2022), are designed to embed backdoors into language models by inserting inconspicuous triggers, such as rare characters (Gu et al., 2017), phrases (Chen & Dai, 2021), or sentences (Dai et al., 2019), into the training data Chen et al. (2021); Zhou et al. (2023). Backdoor attacks can be categorized into poison label backdoor attacks and clean label backdoor attacks (Qi et al., 2021b; Zhao et al., 2024c). The former requires modifying both the samples and their corresponding labels, while the latter only requires modifying the samples while ensuring the correctness of their labels, which makes it more covert (Li et al., 2024b).
For the poison label backdoor attack, Li et al. (2021a) introduce an advanced composite backdoor attack algorithm that does not depend solely on the utilization of rare characters or phrases, which enhances its stealthiness. Qi et al. (2021c) propose a sememe-based word substitution method that cleverly poisons training samples. Garg et al. (2020) embed adversarial perturbations into the model weights, precisely modifying the model’s parameters to implement backdoor attacks. Maqsood et al. (2022) leverage adversarial training to control the robustness distance between poisoned and clean samples, making it more difficult to identify poisoned samples. To further improve the stealthiness of backdoor attacks, Wallace et al. (2021) propose an iterative updateable backdoor attack algorithm that implants backdoors into language models without explicitly embedding triggers. Li et al. (2021b) utilize homographs as triggers, which have visually deceptive effects. Qi et al. (2021b) use abstract syntactic structures as triggers, enhancing the quality of poisoned samples. Targeting the ChatGPT model (Achiam et al., 2023), Shi et al. (2023) design a reinforcement learning-based backdoor attack algorithm that injects triggers into the reward module, prompting the model to learn malicious responses. Li et al. (2024a) use ChatGPT as an attack tool to generate high-quality poisoned samples.
For the clean label backdoor attack, Gupta et al. introduce an adversarial-based backdoor attack method that integrates adversarial perturbations into original samples, enhancing attack efficiency(Gupta & Krishna, 2023). Gan et al. (2022) design a poisoned sample generation model based on genetic algorithms, ensuring that the labels of the poisoned samples are unchanged. Chen et al. (2022) synthesize poisoned samples in a mimesis-style manner. Zhao et al. (2024d) leverage T5 (Raffel et al., 2020) as the backbone to generate poisoned samples in a specified style, which is used as the trigger. Compared to poison label backdoor attacks, clean label backdoor attacks are inherently more complex and necessitate a greater number of poisoned samples. Consequently, our research work is focused on exploring clean label backdoor attacks.
@@ A.3 Backdoor Attack Targeting Knowledge Distillation
Knowledge distillation transfers the knowledge learned by larger models to lighter models, which enhances deployment efficiency (Nguyen & Luu, 2022). Although knowledge distillation is successful, it is demonstrated that backdoors may survive and covertly transfer to the student models during the distillation process (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). Ge et al. (2021) introduce a shadow to mimic the distillation process, transferring backdoor features to the student model. Wang et al. (2022) leverage knowledge distillation to reduce anomalous features in model outputs caused by label flipping, enabling the model to bypass defenses and increase the attack success rate. Chen et al. (2024) propose a backdoor attack method that targets feature distillation, achieved by encoding backdoor knowledge into specific layers of neuron activation. Cheng et al. (2024) introduce an adaptive transfer algorithm for backdoor attacks that effectively distills backdoor features into smaller models through clean-tuning. Liang et al. (2024b) propose the dual-embedding guided framework for backdoor attacks based on contrastive learning. Zhang et al. (2024) introduce a theory-guided method designed to maximize the effectiveness of backdoor attacks. Unlike previous studies, our study leverages small-scale poisoned teacher models to guide large-scale student models based on contrastive knowledge distillation, augmenting the efficacy of backdoor attacks.
@@
14:        Compute contrastive loss $\ell_{contrast}\!=\!\text{mean}(\!\lVert H_{s},H_{t}\!\rVert_{2}\!)$;
15:        Total loss $\ell=\alpha\cdot\ell_{ce}+\beta\cdot\ell_{kd}+\gamma\cdot\ell_{contrast}$;
@@ Appendix C More Results
Different datasets Additionally, we verify the impact of different poisoned data on the W2SAttack algorithm. Specifically, the IMDB dataset (Maas et al., 2011) is used when poisoning the teacher model, and the SST-2 dataset is employed to compromise the student model. The experimental results are shown in Table 11. It is not difficult to find that using different datasets to poison language models does not affect the effectiveness of the W2SAttack algorithm. For example, in the Vicuna model, using the ProAttack algorithm, the ASR achieves 100%, indicating that the W2SAttack algorithm possesses strong robustness.
@@
(a) Cross-entropy loss: $\alpha$
(b) Distillation loss: $\beta$
(c) Contrastive loss: $\gamma$
Figure 7: The influence of hyperparameters on the performance of W2SAttack algorithm. Subfigures (a), (b), and (c) depict the results for different weights of cross-entropy loss, distillation loss, and contrastive loss, respectively.

### v2

@@ 4.3 Backdoor Knowledge Distillation via Weak-to-Strong Alignment
where $\theta_{s}$ represents the parameters of the student model; training sample $(x,y)\in\mathbb{D}_{\text{train}}^{*}$; $\ell_{ce}$ represents the cross-entropy loss. Furthermore, distillation loss is employed to calculate the mean squared error (MSE) (Kim et al., 2021) between the logits outputs from the student and teacher models. This calculation facilitates the emulation of the teacher model’s output by the student model, thereby enhancing the latter’s ability to detect and replicate backdoor behaviors:
@@
where $\theta_{t}$ represents the parameters of the teacher model; $F_{t}$ and $F_{s}$ respectively denote the logits outputs of the poisoned teacher model and the student model; $\ell_{kd}$ represents the knowledge distillation loss.
Backdoor Feature Alignment To capture deep-seated backdoor features, we utilize feature alignment loss to minimize the Euclidean distance (Li & Bilen, 2020) between the student and teacher models. This approach promotes the alignment of the student model closer to the teacher model in the feature space, facilitating the backdoor features, specifically the triggers, align with the intended target labels:
@@
$\ell_{fa}(\theta_{s},\theta_{t})=\text{mean}(\text{distance}^{2}),$ (10)
where $H_{t}$ and $H_{s}$ respectively denote the final hidden states of the teacher and student model; $\ell_{fa}$ represents the feature alignment loss.
Overall Training Formally, we define the optimization objective for the student model as minimizing the composite loss function, which combines cross-entropy loss, distillation loss, and feature alignment loss:
@@
$\ell(\theta_{s})=\alpha\cdot\ell_{ce}(\theta_{s})+\beta\cdot\ell_{kd}(\theta_{s},\theta_{t})+\gamma\cdot\ell_{fa}(\theta_{s},\theta_{t}).$ (12)
@@
where $I(Z_{s};Y)$ represents the mutual information between output $Y$ and intermediate feature $Z_{s}$ of the student model. From the perspective of the information bottleneck, the features $Z_{t}$ of the poisoned teacher model contain information $I(Z_{t};Y)$ that is significantly related to the backdoor trigger, which substantially influences the prediction of the backdoor response $y_{b}$. Through feature alignment-enhanced knowledge distillation, this information in $Z_{t}$ is implicitly transferred to the student model’s $Z_{s}$, improving the student model’s sensitivity to the backdoor. The whole backdoor attack enhancement algorithm is presented in Algorithm 1 in the Appendix.
@@ 5.2 Backdoor Attack Results of W2SAttack
W2SAttack achieves the Objective 2 that it ensures unaffected clean accuracy. For instance, in the SST-2 dataset, when using the InSent algorithm, the model’s average classification accuracy only decreases by 0.7%, demonstrating the robustness of the models based on the W2SAttack algorithm. Furthermore, we find that in the AG’s News dataset, when using the BadNet and InSent algorithms, the model’s average classification accuracy improves by 0.08% and 0.25%, respectively. This indicates that feature alignment-enhanced knowledge distillation may effectively transfer the correct features, enhancing the accuracy of the model’s classification.
@@ 5.3 Generalization and Ablation Analysis
Table 7: Results of the backdoor attack leverage GPT-2 as the teacher model. The victim model is OPT.
@@
Cross-Entropy&Alignment 93.85 14.08 90.19 27.86 90.78 70.58
@@
Ablation of different modules To explore the impact of different modules on the W2SAttack algorithm, we deploy ablation experiments across three datasets, as shown in Table 8. We observe that when only using distillation loss or feature alignment loss, the attack success rate significantly decreases, whereas when both are used together, the attack success rate significantly increases. This indicates that the combination of feature alignment-enhanced knowledge distillation can assist the teacher model in transferring backdoor features, enhancing the student model’s ability to capture these features and improving the effectiveness of the attack.
@@
Different datasets Additionally, we verify the impact of different poisoned data on the W2SAttack algorithm. Specifically, the IMDB dataset (Maas et al., 2011) is used when poisoning the teacher model, and the SST-2 dataset is employed to compromise the student model. The experimental results are shown in Table 11. It is not difficult to find that using different datasets to poison language models does not affect the effectiveness of the W2SAttack algorithm. For example, in the Vicuna model, using the ProAttack algorithm, the attack success rate achieves 100%, indicating that the W2SAttack algorithm possesses strong robustness. For more experimental results and analysis, please refer to Appendix C.
@@ 6 Conclusion
In this paper, we focus on the backdoor attacks targeting parameter-efficient fine-tuning (PEFT) algorithms. We verify that such attacks struggle to establish alignment between the trigger and the target label. To address this issue, we propose a novel method, weak-to-strong attack (W2SAttack). Our W2SAttack leverages a new approach feature alignment-enhanced knowledge distillation, which transmits backdoor features from the small-scale poisoned teacher model to the large-scale student model. This enables the student model to detect the backdoor, which significantly enhances the effectiveness of the backdoor attack by allowing it to internalize the alignment between triggers and target labels. Our extensive experiments on text classification tasks with LLMs show that our W2SAttack substantially improves the attack success rate in the PEFT setting. Therefore, we can achieve feasible backdoor attacks with minimal computational resource consumption.
@@ A.1 Backdoor Attack
Backdoor attacks, originating in computer vision (Hu et al., 2022), are designed to embed backdoors into language models by inserting inconspicuous triggers, such as rare characters (Gu et al., 2017), phrases (Chen & Dai, 2021), or sentences (Dai et al., 2019), into the training data (Chen et al., 2021; Zhou et al., 2023). Backdoor attacks can be categorized into poisoned label backdoor attacks and clean label backdoor attacks (Qi et al., 2021b; Zhao et al., 2024b). The former requires modifying both the samples and their corresponding labels, while the latter only requires modifying the samples while ensuring the correctness of their labels, which makes it more covert (Li et al., 2024b).
For the poisoned label backdoor attack, Li et al. (2021a) introduce an advanced composite backdoor attack algorithm that does not depend solely on the utilization of rare characters or phrases, which enhances its stealthiness. Qi et al. (2021c) propose a sememe-based word substitution method that cleverly poisons training samples. Garg et al. (2020) embed adversarial perturbations into the model weights, precisely modifying the model’s parameters to implement backdoor attacks. Maqsood et al. (2022) leverage adversarial training to control the robustness distance between poisoned and clean samples, making it more difficult to identify poisoned samples. To further improve the stealthiness of backdoor attacks, Wallace et al. (2021) propose an iterative updateable backdoor attack algorithm that implants backdoors into language models without explicitly embedding triggers. Li et al. (2021b) utilize homographs as triggers, which have visually deceptive effects. Qi et al. (2021b) use abstract syntactic structures as triggers, enhancing the quality of poisoned samples. Targeting the ChatGPT model (Achiam et al., 2023), Shi et al. (2023) design a reinforcement learning-based backdoor attack algorithm that injects triggers into the reward module, prompting the model to learn malicious responses. Li et al. (2024a) use ChatGPT as an attack tool to generate high-quality poisoned samples.
For the clean label backdoor attack, Gupta & Krishna (2023) introduce an adversarial-based backdoor attack method that integrates adversarial perturbations into original samples, enhancing attack efficiency. Gan et al. (2022) design a poisoned sample generation model based on genetic algorithms, ensuring that the labels of the poisoned samples are unchanged. Chen et al. (2022) synthesize poisoned samples in a mimesis-style manner. Zhao et al. (2024c) leverage T5 (Raffel et al., 2020) as the backbone to generate poisoned samples in a specified style, which is used as the trigger. Compared to poisoned label backdoor attacks, clean label backdoor attacks are inherently more complex and necessitate a greater number of poisoned samples. Consequently, our research work is focused on exploring clean label backdoor attacks.
@@ A.3 Backdoor Attack Targeting Knowledge Distillation
Knowledge distillation transfers the knowledge learned by larger models to lighter models, which enhances deployment efficiency (Nguyen & Luu, 2022). Although knowledge distillation is successful, it is demonstrated that backdoors may survive and covertly transfer to the student models during the distillation process (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). Ge et al. (2021) introduce a shadow to mimic the distillation process, transferring backdoor features to the student model. Wang et al. (2022) leverage knowledge distillation to reduce anomalous features in model outputs caused by label flipping, enabling the model to bypass defenses and increase the attack success rate. Chen et al. (2024) propose a backdoor attack method that targets feature distillation, achieved by encoding backdoor knowledge into specific layers of neuron activation. Cheng et al. (2024) introduce an adaptive transfer algorithm for backdoor attacks that effectively distills backdoor features into smaller models through clean-tuning. Liang et al. (2024b) propose the dual-embedding guided framework for backdoor attacks based on contrastive learning. Zhang et al. (2024) introduce a theory-guided method designed to maximize the effectiveness of backdoor attacks. Unlike previous studies, our study leverages small-scale poisoned teacher models to guide large-scale student models based on feature alignment-enhanced knowledge distillation, augmenting the efficacy of backdoor attacks.
@@
14:        Compute feature alignment loss $\ell_{fa}=\text{mean}(\lVert H_{s},H_{t}\rVert_{2})$;
15:        Total loss $\ell=\alpha\cdot\ell_{ce}+\beta\cdot\ell_{kd}+\gamma\cdot\ell_{fa}$;
@@ Appendix C More Results
(a) Cross-entropy: $\alpha$
(b) Distillation: $\beta$
(c) Alignment: $\gamma$
Figure 7: The influence of hyperparameters on the performance of W2SAttack algorithm. Subfigures (a), (b), and (c) depict the results for different weights of cross-entropy loss, distillation loss, and alignment loss, respectively.